OpenStudy (anonymous):

the balance of an account earning compounded interest is found using the formula A = p(1 = r)^t, where p is the principal :the amount invested), r is the interest rate, and t is the time in years. what is the account balance after 6 years, if $1,000 is invested at 8%? round your answer to the nearest whole number. A. $1,080 B. $1,587 C. $1,480 D. $1,800 help?

jimthompson5910 (jim_thompson5910):

A = P(1+r)^t A = 1000(1+0.08)^6 A = ???

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

evaluate that to get your answer

OpenStudy (anonymous):

I got B as the answer because I did 1.08*1.08*1.08*1.08*1.08*1.08+1000

OpenStudy (anonymous):

hanks jim and dev